# Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination

### v1

@@ Abstract
Real-time rendering with global illumination is crucial to afford the user realistic experience in virtual environments. We present a learning-based estimator to predict diffuse indirect illumination in screen space, which then is combined with direct illumination to synthesize globally-illuminated high dynamic range (HDR) results. Our approach tackles the challenges of capturing long-range/long-distance indirect illumination when employing neural networks and is generalized to handle complex lighting and scenarios. From the neural network thinking of the solver to the rendering equation, we present a novel network architecture to predict indirect illumination. Our network is equipped with a modified attention mechanism that aggregates global information guided by spacial geometry features, as well as a monochromatic design that encodes each color channel individually. We conducted extensive evaluations, and the experimental results demonstrate our superiority over previous learning-based techniques. Our approach excels at handling complex lighting such as varying-colored lighting and environment lighting. It can successfully capture distant indirect illumination and simulates the interreflections between textured surfaces well (i.e., color bleeding effects); it can also effectively handle new scenes that are not present in the training dataset.
keywords:
global illumination, neural network, real-time rendering
[…]
Our model efficiently synthesizes HDR results with global illumination predicted in around 12 milliseconds. We highlight that our model is well generalized to handle complex dynamic lighting in scenes, such as varying-colored lighting (left) and environment lighting (right), which are new scenes that have not been present in the training dataset.
@@ 1 Introduction
To address these issues, we adopt a learning-based approach to address the issue of global illumination, following the principles of the rendering equation [19]. Our approach include a learnable module that aggregates global features while utilizing geometry supervision. This module incorporates a modified multi-head attention mechanism [41] to calculate the distribution based on geometry features, which guides the global aggregation of spatial features. This non-local process can be analogized to an integral over screen space, where the geometry-aware attention works as a learnable counterpart of the geometry term. Our proposed module overcomes the network bottleneck of the deep stack of convolutional layers, requiring less computation and fewer parameters while better capturing the long-range dependencies. Moreover, it can be applied to other tasks where geometry supervision is available and long-range dependency modeling is preferred.
Second, we present a monochromatic model design, where each RGB channel is independently processed using a shading generator conditioned on shared features from a geometry encoder. The redundant inter-channel interactions (introduced by the interconnections of convolutional filters) can be eliminated, and this relieves the burden of training data and network capacity and enforces the network to learn the accurate color of indirect light. Furthermore, this makes our model immune to color variations, therefore allowing better generalization to novel and complex scenes and dynamic lighting, which are not present in the training data. Since HDR images can provide a much wider range of brightness and contrast for the high sensitivity of the human visual system to enhance the user’s perception of immersion and presence in the immersive virtual environment [7, 27], our model is trained to synthesize HDR global illumination.
Finally, we created a novel synthetic HDR dataset of indoor scenes for network training based on the public indoor layout dataset 3D-Front [11]. Our dataset comes with well-annotated lighting, texture, and geometry features; the surfaces are also assigned with physically-based materials, providing sufficient diversity for geometry features and visual appearances. We design an adversarial training framework to directly optimize our network on HDR data, where the discriminator and the perceptual loss are jointly used for network optimization.
Our model can produce plausible indirect illumination at the real-time performance, handling global effects, accurately predicting indirect shading with good generalization, as shown in Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination. To summarize, our main contributions include:
  • We propose a neural network for indirect illumination prediction that leverages a global feature aggregation module for geometry features to model long-range dependencies, thereby synthesizing realistic HDR images at a reasonable time per frame.
@@ 2.3 Network Conditioning
Inspired by the great success of attention mechanism in modeling global dependencies [41, 46] and motivated by our task formulation, we propose a feature aggregation method that models long-range dependencies guided by geometry features, where the aggregation weights are obtained from a modified multi-head attention mechanism. Different from the prior work on integrating attention to CNN [46], we 1) explicitly use geometry features to guide the aggregation of global features; then 2) extend it into multi-head attention [41], and find it sufficient to equip only bottleneck layers with attention. Both bring a significant reduction of computational overhead, with no observable performance degradation in our experiments.
@@ 2.4 Global illumination prediction
Several recent works use neural networks to predict global illumination given auxiliary shading and geometry features. Nalbach et al. [31] first introduced CNN into global illumination predicting. They use a U-shaped CNN, taking pixel-wise features as inputs and predicting globally illuminated images. Bi et al. [2] train GANs to predict global illumination, and Dai et al. [6] use the perceptual loss for network optimization. Then Xin et al. [44] used a lightweight CNN and predicted the indirect component similar to the task setup in our work. However, Nalbach et al. [31] used global illumination computed in screen space as their ground truth, and Dai et al. excluded textures in their ground truth data generation. Most previous works used LDR data for training or took no special care in training in linear HDR space. Moreover, they generally use stacked convolutional operations for shading learning, limiting the receptive fields to local neighborhoods, and thus is difficult to predict global effects like indirect illumination.
@@
Figure 1: Overview of our model architecture. GCM: geometry-aware conditional modulation; GFA: geometry-aware feature aggregation. The geometry encoder takes in the auxiliary geometry information of each frame as the input and provides geometry conditional features (GCM) and attention weights (GFA) to the generator. The monochromatic shading generator then independently predicts each color channel of the indirect illumination, given lighting and reflectance of the corresponding channel as input.
@@ 3 Our Method
In this section, we first outline the neural network thinking of the solver to rendering equation. Then, we present our network architecture, including the feature aggregation module and the monochromatic design. Finally, we discuss the technical details, practical considerations about performance and optimization, and the adversarial training framework.
@@ 3.1 Task Formulation
We aim to predict indirect lighting given direct lighting and auxiliary geometry features, including surface normal, depth, reflectance, etc. All these inputs can be efficiently computed using a rasterization-based rendering pipeline in real time. To simplify our task, we assume all Lambertian surfaces within the scene, i.e., the BRDF term $f(\mathrm{p},\omega_{i},\omega_{o})$ is a constant for arbitrary $\omega_{i}$ and $\omega_{o}$. We first briefly review the concepts of rendering equation [19] and physically based rendering; by splitting the incident radiance $L_{i}$ into direct and indirect terms, we obtain an alternate version of the surface form rendering equation:
@@
Summarization To predict indirect illumination, the geometry information (surface normals $\mathbf{N}$, depth $\mathbf{D}$, diffuse reflectance $\mathbf{R}$, bold type indicates image data) and direct lighting $\mathbf{L}_{\mathrm{d}}$ are provided as conditioning features. The network is expected to predict the indirect shading component $\mathbf{S}_{\mathrm{ind}}$, from which we can obtain the full indirect illumination $\mathbf{L}_{\mathrm{ind}}$ by multiplying the reflectance, i.e., $\mathbf{L}=\mathbf{R}\cdot\mathbf{S}$. The final global illumination $\mathbf{L}$ is acquired by adding direct illumination $\mathbf{L}_{\mathrm{d}}$ (provided as inputs) to the predicted indirect illumination, i.e., $\mathbf{L}=\mathbf{L}_{\mathrm{d}}+\mathbf{L}_{\mathrm{ind}}$. Note that all the above computations are pixel-wise operations in linear space, thus all the components should be represented with linear values and stored as HDR images.
@@ 3.2 Model Design
Geometry-Aware Feature Aggregation Fully convolutional networks (FCN) will result in a limited-sized receptive field [44, 31, 6], i.e., the output pixel value is conditioned only on a local region. While deep stacks of convolutional operations can enlarge the receptive field, it also brings prohibitive computational costs, making network optimization difficult. In this work, we expect the network to generate indirect illumination from features of all locations within the image space. A learnable module is needed that can efficiently aggregate global features guided by geometry.
@@
where $i$ and $j$ are indices to the features within image space, $Q$ and $K$ are linear transformations (implemented as $1\times 1$ convolutions). The encoded features $\hat{f_{g}}$ are again concatenated with original geometry inputs (down-sampled to the current size) to preserve them from being washed away. The attention distribution of $i$ is normalized for all spatial location $j$ using $\mathrm{softmax}$, and is then used to get the aggregated global feature $\hat{x}^{i}$ for location $i$ as:
@@
where $x^{i}$ is the residual connection applied to outputs, and $V$ is an extra linear transformation applied to the original features while preserving its original dimensionality. Intuitively, this process is similar to the process of integrating over screen space, where the attention mechanism calculates some sort of mutual contribution between two points for shading based on their geometry features (just like a learnable counterpart of the geometry term, Eq. 1). By attending to the projected subspace of geometry features, it learns to determine whether the feature $x_{j}$ of location $j$ is useful and thus should be aggregated into the current location $i$ (see Fig. 2).
@@
Specifically, we interpret the rendering process as a mapping $F:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{X}$ is the high-dimensional inputs encoding the scene geometry, camera, materials, etc.; and $\mathcal{Y}$ is the rendered 2D image, thus rendering can be viewed as evaluating the function $Y=F(X)$. Since the rendering process is wavelength-independent (thus is channel-independent if the RGB color model is used), this function can be further written in a channel-independent way, i.e., $Y^{c}=F^{c}(X^{c})$, where $c\in\{\mathrm{r},\mathrm{g},\mathrm{b}\}$ indicates the specific channel. However, in the general learning-based methods, the mapping between inputs and outputs is written as $\hat{Y}=G(X;\theta)$, where $\theta$ is the model parameter, and $\hat{Y}$ is the predicted indirect illumination. Note that each color channel of the RGB inputs (i.e., reflectance and lighting) contributes to all three channels of $\hat{Y}$ via the interconnections within the convolutional filters, which is redundant and inconsistent with the typical rendering process.
Based on this observation, our monochromatic shading generator processes each color channel independently, i.e., for each input with RGB channels, the shading network parallelly takes in each channel and predicts the shading of that channel. Note that the same shading network is shared between different color channels. We could thus interpret the prediction process as:
@@ 4 Synthetic Dataset
We created a novel synthetic dataset of indoor scenes (see Fig. 4) to train our network, which is based on the 3D-Front [11] public interior dataset with over 60k distinct house layouts. To the best of our knowledge, there is no publicly available dataset containing completely annotated lighting (with factorized shading and reflectance, separated direct/indirect and diffuse/specular components) and geometry features which adapt well to our task. Our dataset will be released publicly online upon publication to fill the absence of a high-quality dataset for learning-based shading prediction tasks.
Our synthetic dataset contains over 30,000 high dynamic range (HDR) images rendered from 1.5k indoor layouts from 3D-Front, each with pixel-wise annotations of lighting, texture, and geometry features. While all textures in 3D-Front are restricted to diffuse materials, and some surfaces have no materials, we also assign physically-based materials to the surfaces, which provide highly diverse visual effects and improve visual realism compared to similar synthetic datasets as [38, 48]. Note that our dataset also applies to many other computer vision tasks, such as indoor scene understanding, inverse rendering, intrinsic decomposition, etc. Please see the supplementary document for more details and illustrations about material assignation and dataset generation.
@@
Texture assignment To achieve more realism on our synthetic dataset, we use an open-source CC0 texture library [8] which includes over 1,300 types of physically based materials covering various surfaces such as wood, metal, glass, etc. We randomly assign certain materials based on the object type in a rule-based style, using a similar strategy with [9]. For example, for each wall surface, we uniformly sample one material among all wood, marble, or concrete-typed materials. This rule-based technique for random texture assignment significantly reduces the effort of manually assigning the textures for each object while preserving sufficient consistency and diversity for indoor layouts.
Light sources placement We randomly sample the placement of light sources. To obtain better approximations of direct lighting at real-time performance, we use directional and point/spherical light sources only. We place point and spherical light sources for each room at random positions near the ceiling. We let all spherical light sources’ radius and intensity vary within a certain range to acquire more diverse visual appearances.
@@ 5.2 Results and Discussions
Table 1: Results of averaged metrics on our test set. Our approach shows superior performance over others. "Ambient" refers to applying a constant, uniform incident radiance to approximate indirect light.
@@
We use three evaluation metrics to compare the performance of different methods, including SSIM, PSNR, and LPIPS [47]. The LPIPS metric calculates perceptual similarity based on high-level features in a pre-trained AlexNet [24], which is more robust to noise and better resembles human choice. We compare two recent works on global illumination prediction, including Nalbach et al. [31] and Xin et al. [44]. Note that Nalbach et al. [31] directly reconstructs the global illumination image, while Xin et al. [44] predicts the indirect component, similar to our task setup. For Nalbach et al.’s implementation, we use the "mono" version they described. We train all models on our dataset since the dataset they used is not publicly available. We also compare our method with the commonly used models in image-to-image translation or image style transfer tasks, including Pix2pix [17] and ResNet [14] with the same training setup and loss function as our model. All evaluation results are obtained on our test set, including 1000 rendered frames from our synthetic dataset and about 200 frames from public scenes we collected, e.g., PinkRoom, LivingRoom, Gallery, and Conference. The averaged results of the evaluation metrics are shown in Tab. 1. We refer readers to the supplemental document for additional visuals on our test set.

### v2

@@ Abstract
Real-time rendering with global illumination is crucial to afford the user realistic experience in virtual environments. We present a learning-based estimator to predict diffuse indirect illumination in screen space, which is then combined with direct illumination to synthesize globally-illuminated high dynamic range (HDR) results. Our approach tackles the challenges of capturing long-range/long-distance indirect illumination when employing neural networks and is generalized to handle complex lighting and scenarios. From the neural network thinking of the solver to the rendering equation, we present a novel network architecture to predict indirect illumination. Our network is equipped with a modified attention mechanism that aggregates global information guided by spatial geometry features, as well as a monochromatic design that encodes each color channel individually. We conducted extensive evaluations, and the experimental results demonstrate our superiority over previous learning-based techniques. Our approach excels at handling complex lighting, such as varying-colored lighting and environment lighting. It can successfully capture distant indirect illumination and simulate the interreflections between textured surfaces well (i.e., color bleeding effects); it can also effectively handle new scenes that are not present in the training dataset.
Figure 1: Our model efficiently synthesizes HDR results with global illumination predicted in around 12 milliseconds. We highlight that our model is well generalized to handle complex dynamic lighting in scenes, such as varying-colored lighting (left) and environment lighting (right), which are new scenes that have not been present in the training dataset.
@@ 1 Introduction
To address these issues, we adopt a learning-based approach to address the issue of global illumination, following the principles of the rendering equation [20]. Our approach includes a learnable module that aggregates global features while utilizing geometry supervision. This module incorporates a modified multi-head attention mechanism [43] to calculate the distribution based on geometry features, which guides the global aggregation of spatial features. This non-local process can be analogized to an integral over screen space, where the geometry-aware attention works as a learnable counterpart of the geometry term. Our proposed module overcomes the network bottleneck of the deep stack of convolutional layers, requiring less computation and fewer parameters while better capturing the long-range dependencies. Moreover, it can be applied to other tasks where geometry supervision is available and long-range dependency modeling is preferred.
Second, we present a monochromatic model design, where each RGB channel is independently processed using a shading generator conditioned on shared features from a geometry encoder. The redundant inter-channel interactions (introduced by the interconnections of convolutional filters) can be eliminated, and this relieves the burden of training data and network capacity, and enforces the network to learn the accurate color of indirect light. Furthermore, this makes our model immune to color variations, therefore allowing better generalization to novel and complex scenes and dynamic lighting, which are not present in the training data. Since HDR images can provide a much wider range of brightness and contrast for the high sensitivity of the human visual system to enhance the user’s perception of immersion and presence in the immersive virtual environment [7, 28], our model is trained to synthesize HDR global illumination.
Finally, we created a novel synthetic HDR dataset of indoor scenes for network training based on the public indoor layout dataset 3D-Front [11]. Our dataset comes with well-annotated lighting, texture, and geometry features; the surfaces are also assigned with physically based materials, providing sufficient diversity for geometry features and visual appearances. We design an adversarial training framework to directly optimize our network on HDR data, where the discriminator and the perceptual loss are jointly used for network optimization.
Our model can produce plausible indirect illumination at real-time performance, handling global effects, accurately predicting indirect shading with good generalization, as shown in Fig. 1. To summarize, our main contributions include:
  • We propose a neural network for indirect illumination prediction that leverages a global feature aggregation module for geometry features to model long-range dependencies, thereby synthesizing realistic HDR images with high performance.
@@ 2.3 Network Conditioning
Inspired by the great success of attention mechanisms in modeling global dependencies [43, 48] and motivated by our task formulation, we propose a feature aggregation method that models long-range dependencies guided by geometry features, where the aggregation weights are obtained from a modified multi-head attention mechanism. Different from the prior work on integrating attention to CNN [48], we 1) explicitly use geometry features to guide the aggregation of global features; then 2) extend it into multi-head attention [43], and find it sufficient to equip only bottleneck layers with attention. Both bring a significant reduction of computational overhead, with no observable performance degradation in our experiments.
@@ 2.4 Global illumination prediction
Several recent works use neural networks to predict global illumination given auxiliary shading and geometry features. Nalbach et al. [32] first introduced CNN into global illumination prediction. They use a U-shaped CNN, taking pixel-wise features as inputs and predicting globally illuminated images. Bi et al. [2] train GANs to predict global illumination, and Dai et al. [6] use the perceptual loss for network optimization. Then Xin et al. [46] used a lightweight CNN and predicted the indirect component similar to the task setup in our work. However, Nalbach et al. [32] used global illumination computed in screen space as their ground truth, and Dai et al. excluded textures in their ground truth data generation. Most prior works used LDR data for training or took no special care in training in linear HDR space. Moreover, they generally use stacked convolutional operations for shading learning, limiting the receptive fields to local neighborhoods, and thus it is difficult to predict global effects like indirect illumination.
@@
Figure 2: Overview of our model architecture. GCM: geometry-aware conditional modulation; GFA: geometry-aware feature aggregation. The geometry encoder takes in the auxiliary geometry information of each frame as input and provides geometry conditional features (GCM) and attention weights (GFA) to the generator. The monochromatic shading generator then independently predicts each color channel of the indirect illumination, given the lighting and reflectance of the corresponding channel as input.
@@ 3 Our Method
We first outline the neural network thinking of the solver for rendering equations. Then, we present our network architecture, including the feature aggregation module and the monochromatic design. Finally, we discuss the technical details, practical considerations about performance and optimization, and the adversarial training framework.
@@ 3.1 Task Formulation
We aim to predict indirect lighting given direct lighting and auxiliary geometry features, including surface normal, depth, reflectance, etc. All these inputs can be efficiently computed using a rasterization-based rendering pipeline in real time. For simplicity, we assume all Lambertian surfaces within the scene, i.e., the BRDF term $f(\mathrm{p},\omega_{i},\omega_{o})$ is a constant for arbitrary $\omega_{i}$ and $\omega_{o}$. We first briefly review the concepts of rendering equation [20] and physically based rendering; by splitting the incident radiance $L_{i}$ into direct and indirect terms, we derive an alternative form of the surface rendering equation:
@@
Summarization To predict indirect illumination, the geometry information (surface normals $\mathbf{N}$, depth $\mathbf{D}$, diffuse reflectance $\mathbf{R}$, bold type indicates image data) and direct lighting $\mathbf{L}_{\mathrm{d}}$ are provided as conditioning features. The network is expected to predict the indirect shading component $\mathbf{S}_{\mathrm{ind}}$, from which we can obtain the full indirect illumination $\mathbf{L}_{\mathrm{ind}}$ by multiplying the reflectance, i.e., $\mathbf{L}=\mathbf{R}\cdot\mathbf{S}$. The final global illumination $\mathbf{L}$ is acquired by adding direct illumination $\mathbf{L}_{\mathrm{d}}$ (provided as inputs) to the predicted indirect illumination, i.e., $\mathbf{L}=\mathbf{L}_{\mathrm{d}}+\mathbf{L}_{\mathrm{ind}}$. Note that all the above computations are pixel-wise operations in linear space; thus, all the components should be represented with linear values and stored as HDR images.
@@ 3.2 Model Design
Geometry-Aware Feature Aggregation Fully convolutional networks (FCN) will result in a limited-sized receptive field [46, 32, 6], i.e., the output pixel value is conditioned only on a local region. While deep stacks of convolutional operations can enlarge the receptive field, they also bring prohibitive computational costs, making network optimization difficult. In this work, we expect the network to generate indirect illumination from features of all locations within the image space. A learnable module is needed that can efficiently aggregate global features guided by geometry.
@@
where $i$ and $j$ are indices to the features within image space, $Q$ and $K$ are linear transformations (implemented as $1\times 1$ convolutions). Specifically, $F$ and $G$ correspond to these two transformations in the geometry encoder, representing the Query ($Q$) and Key ($K$) projections applied to geometry features, as shown in Fig. 3. The encoded features $\hat{f_{g}}$ are again concatenated with original geometry inputs (down-sampled to the current size) to preserve them from being washed away. The attention distribution of $i$ is normalized for all spatial location $j$ using $\mathrm{softmax}$, and is then used to get the aggregated global feature $\hat{x}^{i}$ for location $i$ as:
@@
where $x^{i}$ is the residual connection applied to outputs, and $V$ is an extra linear transformation applied to the original features while preserving their original dimensionality. Intuitively, this process is similar to the process of integrating over screen space, where the attention mechanism calculates some sort of mutual contribution between two points for shading based on their geometry features (just like a learnable counterpart of the geometry term, Eq. 1). By attending to the projected subspace of geometry features, it learns to determine whether the feature $x_{j}$ of location $j$ is useful and thus should be aggregated into the current location $i$ (see Fig. 3).
@@
Specifically, we interpret the rendering process as a mapping $F:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{X}$ is the high-dimensional inputs encoding the scene geometry, camera, materials, etc., and $\mathcal{Y}$ is the rendered 2D image, thus, rendering can be viewed as evaluating the function $Y=F(X)$. Since the rendering process is wavelength-independent (thus is channel-independent if the RGB color model is used), this function can be further written in a channel-independent way, i.e., $Y^{c}=F^{c}(X^{c})$, where $c\in\{\mathrm{r},\mathrm{g},\mathrm{b}\}$ indicates the specific channel. However, in the general learning-based methods, the mapping between inputs and outputs is written as $\hat{Y}=G(X;\theta)$, where $\theta$ is the model parameter, and $\hat{Y}$ is the predicted indirect illumination. Note that each color channel of the RGB inputs (i.e., reflectance and lighting) contributes to all three channels of $\hat{Y}$ via the interconnections within the convolutional filters, which is redundant and inconsistent with the typical rendering process.
Based on this observation, our monochromatic shading generator processes each color channel independently, i.e., for each input with RGB channels, the shading network takes in each channel in parallel and predicts the shading of that channel. Note that the same shading network is shared between different color channels. We could thus interpret the prediction process as:
@@ 4 Synthetic Dataset
We created a novel synthetic dataset of indoor scenes (see Fig. 5) to train our network, which is based on the 3D-Front [11] public interior dataset with over 60k distinct house layouts. To the best of our knowledge, there is no publicly available dataset containing completely annotated lighting (with factorized shading and reflectance, separated direct/indirect and diffuse/specular components) and geometry features that adapt well to our task. Our dataset will be released publicly online upon publication to fill the absence of a high-quality dataset for learning-based shading prediction tasks.
Our synthetic dataset contains over 30,000 high dynamic range (HDR) images rendered from 1.5k indoor layouts from 3D-Front, each with pixel-wise annotations of lighting, texture, and geometry features. While all textures in 3D-Front are restricted to diffuse materials, and some surfaces have no materials, we also assign physically based materials to the surfaces, which provide highly diverse visual effects and improve visual realism compared to similar synthetic datasets as [39, 50]. Note that our dataset also applies to many other computer vision tasks, such as indoor scene understanding, inverse rendering, intrinsic decomposition, etc. Please see the supplementary document for more details and illustrations about material assignment and dataset generation.
@@
Texture assignment To achieve more realism on our synthetic dataset, we use an open-source CC0 texture library [8] which includes over 1,300 types of physically based materials covering various surfaces such as wood, metal, glass, etc. We randomly assign certain materials based on the object type in a rule-based style, using a similar strategy with [9]. For example, for each wall surface, we uniformly sample one material among all wood, marble, or concrete-type materials. This rule-based technique for random texture assignment significantly reduces the effort of manually assigning the textures for each object while preserving sufficient consistency and diversity for indoor layouts.
Light sources placement We randomly sample the placement of light sources. To approximate direct lighting at real-time performance, we use directional and point/spherical light sources only. We place point and spherical light sources for each room at random positions near the ceiling, and let all spherical light sources’ radius and intensity vary within a certain range to acquire more diverse visual appearances.
@@ 5.2 Results and Discussions
Table 1: Validation on our test set demonstrates that our method achieves superior performance compared to existing approaches. "Ambient" refers to applying a constant, uniform incident radiance to approximate indirect light.
@@
We use three evaluation metrics to compare the performance of different methods, including SSIM, PSNR, and LPIPS [49]. The LPIPS metric calculates perceptual similarity based on high-level features in a pre-trained AlexNet [25], which is more robust to noise and better resembles human choice. We compare two recent works on global illumination prediction, including Nalbach et al. [32] and Xin et al. [46]. Our method adopts a neural rendering approach based on shading space, predicting indirect illumination from screen space to achieve real-time HDR global illumination estimation. Its core idea is to efficiently simulate long-range indirect light transport through an image-to-image learning paradigm, without relying on explicit geometric modeling or physical path integration, thus belonging to a fast, screen-space neural rendering framework. In contrast, [51] focuses on 3D explicit scene modeling and object-level light transport representation. It implicitly models the illumination contribution of each object in the scene, superposing them in latent space to solve for global illumination. These methods differ fundamentally in technical routes, input spaces, and problem formulations. Therefore, such scene-space neural rendering methods [51, 41, 14] are not included for direct comparisons.
Note that Nalbach et al. [32] directly reconstructs the global illumination image, while Xin et al. [46] predicts the indirect component, similar to our task setup. For Nalbach et al.’s implementation, we use the "mono" version they described. We train all models on our dataset since the dataset they used is not publicly available. We also compare our method with the commonly used models in image-to-image translation or image style transfer tasks, including Pix2pix [18] and ResNet [15] with the same training setup and loss function as our model. All evaluation results are obtained on our test set, including 1000 rendered frames from our synthetic dataset and about 200 frames from public scenes we collected, e.g., PinkRoom, LivingRoom, Gallery and Conference. The averaged results of the evaluation metrics are shown in Tab. 1.